Doc. Math. J. DMV 648

Section 12. Probability and Statistics

Lattice Point Problems

and the Central Limit Theorem

in Euclidean Spaces

cf. Vol. III, p. 245–255

F. Götze

Errata

1) Replace in formula (1.2) exp\{− \frac{v}{2} \} by Γ( \frac{d}{2} + 1)( \frac{v}{2} ) ^{\frac{d}{2}} exp\{− \frac{v}{2} \}.

2) On the same page replace line 14–16 from bottom by:

... values in the interval [0, n − 1] with probability c > χ(1)/2, n > n_{0}. Thus there exists an integer j such that

P {T_{n} = j n^{−1}} ≥ c n^{−1}.

3) In the display formula after (5.6) replace κ = c_{2}δ^{−4/d} by κ = c_{2}δ^{4/d}.

Friedrich Götze Fakultät für Mathematik Universität Bielefeld Postfach 100131 33501 Bielefeld Germany

goetze@mathematik.Uni-Bielefeld.DE

Documenta Mathematica · Extra Volume ICM 1998